# Optimistic Safety for Linearly-Constrained Online Convex Optimization

### v1

@@ 2.3 Assumptions
0:  Action set $\mathcal{X}$, Horizon $T$, Online Algorithm $\mathcal{A}$, Confidence set radius $\beta_{t}$, Regularization $\lambda$.
@@
5:     Update pessimistic action set: $\mathcal{Y}_{j}^{p}=\left\{x\in\mathcal{X}:\hat{A}_{j}x+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq b\right\}$
6:     Update optimistic action set: $\mathcal{Y}_{j}^{o}:=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq b\right\}$
@@
13:        $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$.
14:        $S_{t+1}=S_{t}+y_{t}x_{t}^{\top}$.
15:        $t=t+1$.
16:     end while
17:     Terminate $\mathcal{A}$.
18:     $j=j+1$.
19:  end while
3 Algorithm
To address the stated problem, we propose the algorithm Optimistically Safe OCO (OSOCO) given in Algorithm 1. The key idea behind this algorithm is that it maintains both an optimistic action set, which overestimates the feasible set, and a pessimistic action set, which underestimates the feasible set, to simultaneously ensure low regret and constraint satisfaction. Since the optimistic set contains the true feasible set and therefore the optimal action, we ensure low regret by running an online optimization algorithm over the optimistic set. We then scale down the action selected by the online optimization algorithm in each round until it is in the pessimistic set and therefore satisfies the constraints. Notably, our approach of performing regret minimization while simultaneously learning the constraints differs from existing approaches (e.g. Chaudhary & Kalathil (2022); Chang et al. (2023)) which first perform pure exploration to learn the constraints and then perform regret minimization.
We give the details on OSOCO in Section 3.1. Then, in Section 3.2, we give the requirements on the online optimization algorithm $\mathcal{A}$ that is instantiated by OSOCO.
@@ 3.1 Details on OSOCO
The OSOCO algorithm operates over a number of phases, indexed by $j$. We describe the initialization of each phase, the update in each round, and then the phase termination criteria in the following sections.
@@
At the beginning of each phase $j$, the regularized least-squares estimator of the constraint $\hat{A}_{j}$ is calculated (line 4), which is then used to construct the pessimistic action set $\mathcal{Y}_{j}^{p}$ (line 5) and the optimistic action set $\mathcal{Y}_{j}^{o}$ (line 6). Then, the online optimization algorithm $\mathcal{A}$ is initialized with the optimistic action set (line 7). To appropriately construct the pessimistic and optimistic action sets, we use the confidence set for least-squares estimators from Abbasi-Yadkori et al. (2011), which is given in the following theorem.
Theorem 3.1 (Theorem 2 in Abbasi-Yadkori et al. (2011)).
Fix some $\delta\in(0,1)$ and let $\check{A}_{t}=S_{t}V_{t}^{-1}$. Then, with probability at least $1-\delta$, it holds for all $t\geq 1$ and $x\in\mathbb{R}^{d}$ that
$(\check{A}_{t}-A)x\in\beta_{t}\|x\|_{V_{t}^{-1}}\mathbb{B}_{\infty}$
where
$\beta_{t}:=\rho\sqrt{d\log\left(\frac{1+(t-1)D^{2}/\lambda}{\delta/n}\right)}.$ (1)
It immediately follows from the validity of this confidence set that the optimistic set (resp. pessimistic set) is a superset (resp. subset) of the true feasible set for all rounds with high probability, i.e. $\mathcal{Y}_{j}^{p}\subseteq\mathcal{Y}\subseteq\mathcal{Y}_{j}^{o}$ for all $j$ with probability at least $1-\delta$.
@@
The phase ends when the determinant of the Gram matrix ($V_{t+1}$) has doubled (line 8), after which it terminates $\mathcal{A}$ (line 17) and starts a new phase. This type of “rare update” ensures that the phase only ends when a significant amount of new constraint information has been collected. As such, it results in a small number $\mathcal{O}(d\log(T))$ of phases and therefore there will only be a small cost to restarting the OCO algorithm in each phase. Note that a related form of rare updating was used as early as Abbasi-Yadkori et al. (2011) to reduce the number of determinant calculations in stochastic linear bandits.
@@ 3.2 Requirements of Online Algorithm $\mathcal{A}$
Definition 3.2 (Regret Bound on $\mathcal{A}$).
Consider an online optimization setting where in each round $t$, the player chooses $x_{t}$ from the action set $\mathcal{X}$, suffers the cost $f_{t}(x_{t})$ and then observes $f_{t}$. Importantly,
  • $\mathcal{X}$ is not necessarily convex,
  • the game can be terminated at any time (and therefore the final round $T$ is not known to the player).
Also, assume that all $f_{t}$ are convex and that Assumptions 2.1, 2.2 and 2.6 hold. We further assume that $\mathcal{A}$ chooses a distribution $p_{t}$ in each round, which it uses to sample $x_{t}$. [^3] We assume that $\mathcal{A}$ is randomized because an online optimization algorithm needs randomness to achieve sublinear regret when the action set is not convex, see Cesa-Bianchi & Lugosi (2006). Then, $C_{\mathcal{A}}$ is defined as a quantity such that
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x_{t})\leq C_{\mathcal{A}}\sqrt{T},$
@@
Most algorithms that admit such a regret bound are those designed for the online nonconvex optimization setting as they can handle the nonconvex action set (and then the “doubling trick” can be used to handle the unknown horizon). Some examples of applicable algorithms are FTPL Kalai & Vempala (2005); Agarwal et al. (2019); Suggala & Netrapalli (2020) and continuous versions of the Hedge algorithm Freund & Schapire (1997); Maillard & Munos (2010); Krichene et al. (2015) which enjoy $\mathcal{O}(d^{3/2}\sqrt{T})$ and $\tilde{\mathcal{O}}(\sqrt{dT})$ regret, respectively. We explicitly give several example algorithms and their regret guarantees in Appendix D. The primary drawback of these algorithms is that they are often computationally inefficient. In Section 6, we provide a more computationally efficient version of OSOCO at the cost of a slightly worse regret bound $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
@@ 4 Analysis
In this section, we provide regret and safety guarantees for OSOCO. In particular, Theorem 4.1 shows that, with high probability, OSOCO ensures the constraints are always satisfied and that the regret is $\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$ where $C_{\mathcal{A}}$ is the regret constant of the OCO algorithm $\mathcal{A}$ that is used by OSOCO (see Definition 3.2). It immediately follows that the regret of OSOCO is $\tilde{\mathcal{O}}(d\sqrt{T})$ when $C_{\mathcal{A}}$ is $\tilde{\mathcal{O}}(\sqrt{d})$, which holds when algorithms such as Hedge Freund & Schapire (1997) are used for $\mathcal{A}$ (as given in Appendix D).
Theorem 4.1.
Let $C_{\mathcal{A}}$ be the regret constant of $\mathcal{A}$ as defined in Definition 3.2. Fix some $\delta\in(0,1/2)$ and choose $\lambda=\max(1,D^{2})$, and $\beta_{t}$ according to Theorem 3.1. Then if $T\geq 3$, with probability at least $1-2\delta$, the regret of OSOCO (Algorithm 1) satisfies
and it holds that $Ax_{t}\leq b$ for all $t\in[T]$.
The proof of Theorem 4.1 (given in Appendix A) relies on a decomposition of the regret as
$R_{T}$ $=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\underbrace{\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})}_{\text{Cost of Safety}}+\underbrace{\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})}_{\text{Optimistic Regret}}$
We describe the analysis for the cost of safety in Section 4.1 and the optimistic regret in Section 4.2.
4.1 Cost of Safety
In this section, we study the cost that OSOCO incurs by ensuring constraint satisfaction, which we call the cost of safety. Specifically, the following lemma bounds the cumulative difference in cost between the played actions $x_{t}$ and the optimistic actions $\tilde{x}_{t}$. This is precisely the cost incurred by the algorithm scaling the optimistic action into the pessimistic set (as done in lines 10-11 in Algorithm 1).
Lemma 4.2.
Consider the same setup as Theorem 4.1. Then, it holds that
$\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})\leq 4\frac{DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}.$ (2)
Proof Sketch
The first step in the proof of Lemma 4.2 is to lower bound $\gamma_{t}$. Specifically, given any point $x\in\mathcal{Y}_{j}^{o}$, we find a scaling $\alpha\in[0,1]$ such that the scaled point $\alpha x$ is in $\mathcal{Y}_{j}^{p}$. Using this approach, it can be shown (Lemma A.2) that
$\gamma_{t}\geq 1-\frac{2}{b_{\mathrm{min}}}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$
where $j_{t}$ is the phase that round $t$ belongs to. Then, using the fact that all $f_{t}$ have bounded gradients, the lefthand side of (2) can be bounded with a constant multiple of $\sum_{t=1}^{T}(1-\gamma_{t})$ and therefore with a constant multiple of $\sum_{t=1}^{T}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}}$. From the phase termination criteria (line 17), the determinant of $V_{t}$ is no more than twice as large as the determinant of $\bar{V}_{j_{t}}$, so $\|x_{t}\|_{\bar{V}_{j_{t}}}^{-1}\leq 2\|x_{t}\|_{V_{t}}^{-1}$ (where we implicitly used Lemma 12 in Abbasi-Yadkori et al. (2011)). Also, $\beta_{t}$ is monotone in $t$ so it holds that $\beta_{T}\geq\bar{\beta}_{j_{t}}$ Putting these together, we get a bound in terms of $\beta_{T}\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}$ and therefore we can apply Cauchy-Schwarz and the elliptic potential lemma Dani et al. (2008); Abbasi-Yadkori et al. (2011) to get the result.
4.2 Optimistic Regret
In this section, we study the regret incurred by the optimistic actions, which we call the optimistic regret. This measure of regret is “optimistic” in that it is the regret that the algorithm would incur by using an optimistic estimate of the constraint.
Lemma 4.3.
Consider the same setup as Theorem 4.1. Then, it holds that
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})\leq C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
with probability at least $1-2\delta$.
Proof Sketch
Let $N$ be the number of phases, $t_{j}$ be the first round in the $j$th phase (where $t_{N+1}:=T+1$), and $x^{\star}_{j}$ be an optimal action in phase $j$, i.e.
$x^{\star}_{j}\in\operatorname*{arg\,min}_{x\in\mathcal{Y}_{j}^{o}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)$
It follows from Theorem 3.1 that $\mathcal{Y}$ is a subset of $\mathcal{Y}_{j}^{o}$ for all $j$ with probability at least $1-\delta$. As such, it holds with the same probability that $\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})\leq\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star})$, and therefore that
$\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star})$
$\leq\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})$
$\leq\sum_{j=1}^{N}C_{\mathcal{A}}\sqrt{t_{j+1}-t_{j}}$
$\leq C_{\mathcal{A}}\sqrt{NT},$
where the second inequality is the regret bound of $\mathcal{A}$ and the third inequality is Cauchy-Schwarz. Then, by analyzing the growth of the determinant of $V_{t}$, it can be shown that $N\leq 2d\log(T)$ (Lemma A.7). Finally, applying the Azuma inequality and union bound gives the result.
5 Zero Expected Constraint Violation
Thus far, we have viewed the constraint as being a fixed linear function ($Ax$) that the player gets noisy feedback on ($y_{t}$) and must satisfy at all rounds (at least with high probability). In this section, we take the alternative viewpoint that the constraint feedback $y_{t}$ is a stochastic function that must always be satisfied in expectation. We study this setting in Proposition 5.1 (proven in Appendix B), where we modify the OSOCO algorithm by slightly tightening the constraint, which ensures that $\mathbb{E}[y_{t}]\leq 0$ for all $t$ while only adding constants to the regret bound.
This slight modification of the problem setting allows for comparison with the literature on OCO under time-varying constraints, e.g. Neely & Yu (2017); Cao et al. (2018); Yi et al. (2021); Guo et al. (2022), and, in particular, OCO with stochastic constraints Yu et al. (2017); Castiglioni et al. (2022). Specifically, algorithms were given by Yu et al. (2017), Wei et al. (2020) and Castiglioni et al. (2022) showing that it is possible to guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ expected regret and $\tilde{\mathcal{O}}(\sqrt{T})$ expected violation when the constraints are iid and bounded convex functions, and the player receives full-information feedback on the constraints. We contribute to this literature by showing that it is possible to guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ expected regret and zero expected constraint violation when the constraint functions are linear in expectation and conditionally subgaussian, and the player receives bandit feedback on the constraints.
Proposition 5.1.
Consider the same setup as Theorem 4.1 except with OSOCO modified such that $b\leftarrow b-\mu\mathbf{1}$ for some $\mu\in[0,b_{\mathrm{min}})$. Then, it holds that $\mathbb{E}[y_{t}]\leq b$ for all $t\in[T]$ and
@@ 5 Zero Expected Constraint Violation
$\mathbb{E}[R_{T}]\leq\frac{8DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}+GD+\frac{b_{\mathrm{min}}G}{S}$
when $\mu=b_{\mathrm{min}}/T$ and $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$.

### v3

@@ 3 Algorithm
0:  Action set $\mathcal{X}$, Horizon $T$, Online algorithm $\mathcal{A}$, Confidence set radius $\beta_{t}>0$, Regularization $\lambda>0$, Tightening $\kappa\in[0,b_{\mathrm{min}})$.
@@
5:     Update pessimistic action set:$\mathcal{Y}_{j}^{p}:=\left\{x\in\mathcal{X}:\hat{A}_{j}x+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq b-\kappa\mathbf{1}\right\}$
6:     Update optimistic action set:$\mathcal{Y}_{j}^{o}:=\left\{x\in\mathcal{X}:\hat{A}_{j}x-\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq b-\kappa\mathbf{1}\right\}$
@@
13:        $V_{t+1}=V_{t}+x_{t}x_{t}^{\top},S_{t+1}=S_{t}+y_{t}x_{t}^{\top}$.
14:        $t=t+1$.
15:     end while
16:     Terminate $\mathcal{A}$.
17:     $j=j+1$.
18:  end while
To address the stated problem, we propose the algorithm Optimistically Safe OCO (OSOCO) given in Algorithm 1. This algorithm operates over phases and invokes a generic online optimization algorithm $\mathcal{A}$ in each phase. We give the details on OSOCO in Section 3.1. Then, in Section 3.2, we give the requirements on the online optimization algorithm $\mathcal{A}$.
@@ 3.1 Details on OSOCO
The OSOCO algorithm operates over a number of phases. We describe the initialization of each phase, the update in each round, and then the phase termination criteria in the following.
@@
At the beginning of each phase $j$, the regularized least-squares estimator of the constraint $\hat{A}_{j}$ is calculated (line 4), which is then used to construct the pessimistic action set $\mathcal{Y}_{j}^{p}$ (line 5) and the optimistic action set $\mathcal{Y}_{j}^{o}$ (line 6). Then, the online optimization algorithm $\mathcal{A}$ is initialized with the optimistic action set (line 7). To appropriately construct the pessimistic and optimistic action sets, we use bounds on the constraint function of the form,
$\hat{A}_{j}x-\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}\leq Ax\leq\hat{A}_{j}x+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\mathbf{1}$ (1)
which were shown in Abbasi-Yadkori et al. (2011) to hold for all rounds with high probability given an appropriate choice of confidence radius $\bar{\beta}_{j}$. It is natural that this bound is proportional to the weighted norm $\|x\|_{\bar{V}_{j}^{-1}}$ because this results in the bound being tighter when $x$ is in a direction that has been played more frequently and therefore in a direction for which the least-squares estimator is more accurate. Additionally, we incorporate a tightening parameter $\kappa$ in the optimistic and pessimistic sets which is used to ensure no violation in expectation. Therefore, the bounds on the constraint function in (1) ensure that the optimistic set (resp. pessimistic set) is a superset (resp. subset) of the tightened feasible set $\tilde{\mathcal{Y}}=\{x\in\mathcal{X}:Ax\leq b-\kappa\mathbf{1}\}$ for all phases with high probability, i.e. $\mathcal{Y}_{j}^{p}\subseteq\tilde{\mathcal{Y}}\subseteq\mathcal{Y}_{j}^{o}$ for all $j$ with high probability.
@@
The phase ends when the determinant of the Gram matrix ($V_{t+1}$) has doubled (line 8), after which it terminates $\mathcal{A}$ (line 16) and starts a new phase. This type of “rare update” ensures that the phase only ends when a significant amount of new constraint information has been collected. As such, it results in a small number $\mathcal{O}(d\log(T))$ of phases and therefore there will only be a small cost to restarting the OCO algorithm in each phase. Note that a related form of rare updating was used as early as Abbasi-Yadkori et al. (2011) to reduce the number of determinant calculations in stochastic linear bandits.
@@ 3.2 Requirements of Online Algorithm $\mathcal{A}$
Definition 1 (Regret Bound on $\mathcal{A}$).
Consider an OCO setting where the action set $\mathcal{X}$ is not necessarily convex, and the problem can be terminated at any time (and therefore the final round $T$ is not known to the player). Also, we assume Assumptions 1 and 2 hold, and that $\mathcal{A}$ chooses a distribution $p_{t}$ in each round from which it samples $x_{t}$. [^4] We assume that $\mathcal{A}$ is randomized because an online optimization algorithm needs randomness to achieve sublinear regret when the action set is not convex, see Cesa-Bianchi and Lugosi (2006). Then, $C_{\mathcal{A}}$ is a positive real such that,
$\sum_{t\in[T]}\mathbb{E}_{x_{t}\sim p_{t}}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t\in[T]}f_{t}(x)\leq C_{\mathcal{A}}\sqrt{T},$
@@
Most algorithms that admit such a regret bound are those designed for the online nonconvex optimization setting as they can handle the nonconvex action set (and then the “doubling trick” can be used to handle the unknown horizon). Some examples of applicable algorithms are FTPL (Kalai and Vempala, 2005) and continuous versions of the Hedge algorithm (Freund and Schapire, 1997) which enjoy $\mathcal{O}(d^{3/2}\sqrt{T})$ and $\tilde{\mathcal{O}}(\sqrt{dT})$ regret, respectively. We explicitly give several example algorithms and their regret guarantees in Appendix F. The primary drawback of these algorithms is that they are often computationally inefficient. In Section 5, we provide a more computationally efficient version of OSOCO at the cost of a slightly worse regret bound (i.e. $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ instead of $\tilde{\mathcal{O}}(d\sqrt{T})$).
@@ 4 Analysis
In this section, we provide regret and safety guarantees for OSOCO. The high-probability guarantees and expectation guarantees are given in Sections 4.1 and 4.2, respectively.
4.1 High-probability Guarantees
@@
The following theorem shows that, with high probability, OSOCO ensures no violation $Ax_{t}\leq b$ and that the regret is $\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$, where $C_{\mathcal{A}}$ is the regret constant of the online algorithm $\mathcal{A}$ that is used by OSOCO (see Definition 1). It immediately follows that the regret of OSOCO is $\tilde{\mathcal{O}}(d\sqrt{T})$ when $C_{\mathcal{A}}$ is $\tilde{\mathcal{O}}(\sqrt{d})$, which holds when algorithms such as Hedge (Freund and Schapire, 1997) are used for $\mathcal{A}$ (given in Appendix F).
Theorem 1.
Fix some $\delta\in(0,1/2)$, and set $\lambda=\max(1,D^{2})$, $\beta_{t}=\rho\sqrt{d\log\left(\frac{1+(t-1)D^{2}/\lambda}{\delta/n}\right)}+\sqrt{\lambda}S$ and $\kappa=0$. Then if $T\geq 3$, with probability at least $1-2\delta$, the actions of OSOCO (Algorithm 1) satisfy,
@@
$\begin{split}R_{T}\leq&\mbox{ }4\frac{DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}\\ &+2DG\sqrt{2T\log(1/\delta)}\end{split}$
and $Ax_{t}\leq b$ for all $t\in[T]$.
The proof of Theorem 1 (given in Appendix B), decomposes the regret as,
$R_{T}=\underbrace{\sum_{t\in[T]}(f_{t}(x_{t})-f_{t}(\tilde{x}_{t}))}_{\textrm{Term~{}I}}+\underbrace{\sum_{t\in[T]}(f_{t}(\tilde{x}_{t})-f_{t}(x^{\star})).}_{\textrm{Term~{}II}}$
Intuitively, Term I is the cost of safety, i.e. the cost that the algorithm incurs by maintaining no violation, while Term II is the optimistic regret, i.e. the regret that the algorithm would incur if it played the optimistic actions $\tilde{x}_{t}$ (although this would violate the constraints). We bound each of these in the following lemmas.
Lemma 1 (Cost of Safety).
Under the same conditions as Theorem 1, it holds that
${\textstyle\mathrm{Term\ I}\leq\frac{4DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}.}$ (2)
The key idea behind the proof of Lemma 1 is that, if we can relate Term I to the estimation error of the constraint given in (1), then we can show that it is $\tilde{\mathcal{O}}(\sqrt{T})$ by the elliptic potential lemma (Dani et al. (2008)). One difficulty in doing so is relating the distance between $\tilde{x}_{t}$ and $x_{t}$ to the estimation error. We do so by first establishing a lower bound on the safe scaling $\gamma_{t}$ (line 10) in terms of the estimation error at the optimistic action $\bar{\beta}_{j}\|\tilde{x}_{t}\|_{V_{j}^{-1}}$ and then relating this to the estimation error at the played actions $\bar{\beta}_{j}\|x_{t}\|_{V_{j}^{-1}}$ using the fact that $\tilde{x}_{t}$ and $x_{t}$ are scalar multiples of each other. Then, since all $f_{t}$ have bounded gradients and thus Lipschitz, we can bound Term I in terms of $\|\tilde{x}_{t}-x_{t}\|$ and therefore in terms of $1-\gamma_{t}$ and then $\bar{\beta}_{j}\|x_{t}\|_{V_{j}^{-1}}$. Finally, we adapt the rare updating analysis from Abbasi-Yadkori et al. (2011) to account for the fact that the optimistic and pessimistic sets are updated only at the beginning of each phase, and then apply the elliptic potential lemma to get $\tilde{\mathcal{O}}(\sqrt{T})$ bound.
Lemma 2 (Optimistic Regret).
Under the same conditions as Theorem 1, it holds with probability at least $1-2\delta$ that
$\mathrm{Term\ II}\leq C_{\mathcal{A}}\sqrt{4dT\log(T)}+DG\sqrt{2T\log(1/\delta)}.$
The key idea behind the proof of Lemma 2 is that the optimistic set $\mathcal{Y}_{j}^{o}$ contains the feasible set $\mathcal{Y}$ with high probability (when $\kappa=0$), and therefore we can bound Term II in terms of the sum of the regrets that $\mathcal{A}$ incurs in each phase. Concretely, with $\tau_{j}$ denoting the duration of the $j$th phase and $N$ the number of phases, it holds with high probability,
$\begin{split}&{\textstyle\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})}\\ &{\textstyle\leq\sum_{j=1}^{N}C_{\mathcal{A}}\sqrt{\tau_{j}}\leq C_{\mathcal{A}}\sqrt{NT}\leq C_{\mathcal{A}}\sqrt{4dT\log(T)}},\end{split}$ (3)
where the second inequality is Cauchy-Schwarz and the third inequality uses the fact that $N\leq 2d\log(T)$ thanks to the phase termination criteria. Lastly, we use the Azuma inequality to convert (3) to a high probability bound on Term II.
4.2 Expectation Guarantees
The following proposition shows that OSOCO ensures no expected violation $\mathbb{E}[Ax_{t}]\leq b$ and that the expected regret is $\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$.
Proposition 1.
Choose $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$ and $\kappa=b_{\mathrm{min}}/T$, and set $\lambda$ and $\beta_{t}$ the same as in Theorem 1. Then, if $T\geq 3$, it holds that $\mathbb{E}[Ax_{t}]\leq b$ for all $t\in[T]$ and
$\begin{split}\mathbb{E}[R_{T}]\leq&\mbox{ }\frac{6DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}\\ &+2DG\sqrt{2T\log(1/\delta)}+GD+\frac{b_{\mathrm{min}}G}{S}.\end{split}$
The complete proof of Proposition 1 is given in Appendix C. The key idea behind this proof is that, by tightening the constraint, we can cancel out the violation that might occur if the high probability event in Theorem 1 does not hold. To see this, let $\mathcal{E}$ be the high probability event in Theorem 1 and note that for all $i\in[n]$ and $t\in[T]$,
$\mathbb{E}[a_{i}^{\top}x_{t}]$ $=\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}\}]+\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}^{c}\}]$
$\leq b_{i}-\kappa+2SD\delta\leq b_{i},$
where the first inequality uses that $a_{i}^{\top}x_{t}\leq b_{i}-\kappa$ under $\mathcal{E}$, that $a_{i}^{\top}x_{t}\leq SD$ always and that $\mathbb{P}(\mathcal{E}^{c})\leq 2\delta$, and the last inequality uses the fact that $2SD\delta\leq\kappa$. Then, to characterize the impact that tightening the constraint has on the regret, we bound the distance between the optimal action $x^{\star}$ and the “tightened” feasible set $\tilde{\mathcal{Y}}=\{x\in\mathcal{X}:Ax\leq b-\kappa\mathbf{1}\}$. Specifically, we show that this distance is smaller than $\kappa D/b_{\mathrm{min}}$, and therefore the algorithm only incurs a cost of $\mathcal{O}(1/T)$ per round, or $\mathcal{O}(1)$ in total, for tightening the constraint.
Figure 1: Implementation of HedgeDescent in OSOCO. OSOCO calls HedgeDescent in each phase and passes the family of optimistic sets $\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$ to it. HedgeDescent uses an online gradient descent (OGD) “expert” for each of these sets and then uses Hedge to adaptively select the OGD expert in each round.